OpenStudy (august899):

The arc intercepted by a certain central angle on the unit circle has a length of 11pi/6 units. Find the lengths of the arcs intercepted by the same angle on circles of radius 2 units,3 units,and 4 units. Explain?

jimthompson5910 (jim_thompson5910):

Rule: arc length = (radius)*(angle in radians) if the radius is 1, then the arc length is exactly equal to the angle in radians

OpenStudy (august899):

So since the radius is 2,3,and 4 I just do 2*11pi/6?

jimthompson5910 (jim_thompson5910):

if radius is 2, then yeah 2*(11pi/6) to get... 2*(11pi/6) = (2*11pi)/6 = 22pi/6 = 11pi/3

OpenStudy (august899):

and I do the same for the 3&4?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (august899):

Okay.Thanks for the help!

OpenStudy (august899):

One last question,When the answer came out as 11pi/3 is that the length of the arc or does is it supposed to be reduced?

jimthompson5910 (jim_thompson5910):

11pi/3 is as reduced as it gets

jimthompson5910 (jim_thompson5910):

11pi/3 is the length of the arc if the radius is 2 and the angle is 11pi/6 radians